(2) (a) What is product differentiation? In what market structure is this an essential feature? Briefly explain, at least, three different ways firms in this industry differentiate their products from those of other firms.

                                                                                                                                                                                                        

(b)Suppose a firm in this industry faces the following demand and total cost functions: 

                                                                                             P = 5900 – 10Q  

                                                                                                          3     2

                                                                                            TC = 2Q - 4Q +140Q + 845.

Determine the profit maximizing output the firm should produce, the market clearing price and the total profit (if any) to be earned.

1
Expert's answer
2021-12-19T18:05:37-0500

a.

Product differentiation is the process of separating a product from its competitors in order to entice target consumers to purchase it, hence boosting the firm's revenue. We also mean that the products are distinguishable from the firm's own products when we say differentiation.


Because there are so many enterprises in a monopolistic market, product differentiation is the most important part of competition.


Differentiating products can be done in a variety of ways, including:


Packaging that stands out

In terms of features, there is a distinction to be made. This is the most often used procedure, and it is also the most effective.

Differentiating products based on their size

Differentiating the product's branding in a way that directly challenges the competition

b)

"P = 5900 - 10Q"

Total revenues "= PQ = 5900Q -10Q^\u00b2"

So, Marginal revenue "= 5900 - 20Q"

Total costs "TC = 2Q^\u00b3 - 4Q^\u00b2 + 140Q +845"

So, marginal costs "= 6Q^\u00b2 - 8Q + 140"

At profit maximization, marginal revenue = marginal costs

"=> 5900 - 20Q = 6Q^² - 8Q + 140\\

=> 6Q^² + 12Q - 5760 = 0"

Solving, we get positive value of Q = 30 which is the profit maximizing output.

"P = 5900 - 10Q = 5600"

Profits = Total revenue - Total cost

"= PQ - (2Q^³ - 4Q^² + 140Q + 845)\\

= (5600×30) - (2×30^³ - 4×30^² + 140×30 + 845)\\

= 112,555"